OpenStudy (spicytaco1):

What is the value of u + v + w when u = –10, v = 4, and w = 8? A. –22 B. 2 C. 14 D. 22

OpenStudy (anonymous):

u + v + w when u = –10, v = 4, and w = 8 -10 + 4 + 8 = 2